OpenStudy (anonymous):

What is the limit of (tan^3(2x) )/(x^3) as x approaches 0?

OpenStudy (zarkon):

\[\frac{\tan^3(2x)}{x^3}=\left(\frac{\tan(2x)}{x}\right)^3\]

OpenStudy (zarkon):

\[\lim_{x\to 0}\frac{\tan^3(2x)}{x^3}=\left(\lim_{x\to 0}\frac{\tan(2x)}{x}\right)^3\]

OpenStudy (anonymous):

does that simplify down to sinx/x times 1/cosx?

OpenStudy (zarkon):

\[\tan(2x)=\frac{\sin(2x)}{\cos(2x)}\]

OpenStudy (anonymous):

and then I multiplied by 2/2

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

oh ok tysm :)